Analyze and compare Neiman Marcus and Kohl’s Neiman Marcus Group (NMG) is one of the largest luxury fashion retailers in the world. Kohl’s Corporation (KSS) sells moderately priced private and national branded products through more than 1,100 department stores located throughout the United States. The current assets and current liabilities at the end of a recent year for both companies are as follows (in millions):

Neiman Marcus Kohl’s
Current assets:
Cash $62 $1,074
Inventories 1,125 3,795
Other current assets 147 378
Total current assets $1,334 $5,247
Current liabilities:
Accounts payable $318 $1,507
Other current liabilities 522 1,467
Total current liabilities $840 $2,974

b. Compute the quick ratio for both companies. Round to one decimal place.

Neiman Marcus:

Kohl’s:
Answer #1

Quick ratio = quick assets/current liabilities

Here,

quick assets = cash + other current assets

current liabilities = Total current liabilities

{note: quick assets doesnot include inventories}

Now

Neiman Marcus:

Quick ratio = (62 + 147)/840

= 209/840

= 0.248

= 0.2

Kohl’s:

Quick ratio = (1074 + 378)/2974

= 1452/2974

= 0.48

= 0.5